OpenStudy (anonymous):

reduction formula ∫sec^4x dx

OpenStudy (anonymous):

\[\int\limits \left( \sec ^{2}x \right)\left( \sec ^{2}x \right) dx=\int\limits \left( 1+\tan ^{2}x \right)\sec ^{2}x~dx\] now you can solve.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

the answer is tanx+tan^3x/3+c?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

correct,but it looks you have written x/3 \[\frac{ \tan ^{3}x }{ 3 }\]

OpenStudy (anonymous):

sorry :)